a student on roller skates glides across a floor with a lunch tray. Total mass is 1.3 kg at waist level 1.1 above the floor from the cash register to her table 23.4 m away how much work is done


1
Expert's answer
2021-12-14T13:33:24-0500

"W=F\cdot s\cdot\cos\alpha"


Since the total force acting on the tray is 0 then "W=0" . Answer